# InvisibleInk: High-Utility and Low-Cost Text Generation with Differential Privacy

### v1

@@ D.6 Confidence Intervals
where $\bar{x}$ is the sample mean, $s$ is the sample standard deviation and $m$ is the number of samples. Here, $Z_{\alpha/2}$ is the $(1-\alpha/2)$ percentile of the standard normal distribution. (These are Wald confidence intervals, given under the assumption that error is unbiased and normally distributed.) For a $99\%$ confidence interval, we have $Z_{\alpha/2}=2.576$.
@@ D.7 Computational Resources
Hardware. All experiments requiring a GPU (LLM inferences, $\Delta$PPL calculation, MedNER count calculation etc.) were conducted on one of two machines: (a) with 4 NVIDIA RTX L40S GPUs (with 48G memory each), or (b) with 4 NVIDIA H100 GPUs (with 80GB memory each). Each task used only one GPU at a time. All wall-clock time measurements, if any, are performed using the L40S GPUs. The non-GPU jobs, such as computing MAUVE, were run on a machine with 240 AMD EPYC 9554 CPUs (clock speed: 3.10 GHz) with 64 virtual cores each and total memory of 845GB.
Evaluation Time. The approximate wall clock runtimes for synthetic text generation using various methods are given below. The runtime varies with the size of the model and batch size used, in particular with the maximum number of parallel LLM inferences that may be run on the GPU. For our experiments using TinyLLaMA, a maximum of $16$ inferences were run in parallel on 1 GPU when $B\geq 15$; otherwise, all $B+1$ inferences-per-batch were run in parallel. The runtimes also scale (approximately) linearly with the length of individual generated samples and the number of samples generated. All numbers reported below are for $1000$ generations of $500$ token synthetic samples using MIMIC as the reference dataset.
@@
  • AugPE [36]: $\approx 18$ hours for $7$ variations per iteration and $10$ iterations
@@ Appendix E Additional Experimental Results
  • A comparison of InvisibleInk and AugPE [36] on the MIMIC dataset.
@@
  • Additional experiments designed to highlight the effect of sampling temperature $\tau$ and clipping norm $C$ on the quality of generation.
@@ E.1 Comparison with AugPE
Fig. 7 gives additional experimental results comparing the utility vs computational cost of InvisibleInk and AugPE for various privacy budgets. The trends observed agree with those in Fig. 5; see §5.
Figure 7: Utility vs Computational Cost plots for InvisibleInk and AugPE for varying privacy budgets $\varepsilon$ for 1000 synthetic texts generated for the MIMIC dataset. We compare utility using a variety of metrics — InvisibleInk outperforms AugPE across all settings and evaluation metrics. Wall-clock run time is used to measure the computational cost. We report results for $B+1=4,8,16$ next-token inference calls per generated token and $T_{\mathsf{AugPE}}=1,3,5$ for InvisibleInk and AugPE, respectively.
@@
Table 11: Random sample generation and Variation prompts for MIMIC adapted in the style of AugPE; The candidate synthetic generations at the current iteration are inserted in the space denoted by “<insert candidate synthetic text>” and information about the style in which the LLM should rephrase the text is inserted in the space denoted by “<insert rephrase style>”. System prompt used for both random generation and variation is also given.
@@
The number of variations of the candidate synthetic texts generated in every iteration of AugPE (the parameter $L$ in [36, Alg 1]) is fixed to be $7$. We also limit the total number of sensitive references the algorithm sees to $7\times n$ (where $n$ is the total number of synthetic texts generated). The total number of full-sequence LLM generations per synthetic text is given by $8\times T_{\mathsf{AugPE}}$, where $T_{\mathsf{AugPE}}$ is the number of iterations, and varies from $8-80$. In contrast, we disadvantage our model by limiting the number of effective full-sequence LLM generations to $B+1$, ie, $4-32$. Further, we use the official implementation of AugPE released by Xie et al. [36] [^14] Available at https://github.com/AI-secure/aug-pe and adapt it for generating synthetic samples from MIMIC with minimal changes. In particular, we preserve the same style of random sample generation and text variation prompts used for other datasets; see Tab. 11.
E.2 Top-$k$ Selection
@@
Some additional statistics of selection from the expanded vocabulary described by the Top-$k+$ step of InvisibleInk are presented in Tab. 12. In particular, we show the following:
  • Effective $k$: the size of the vocabulary $V_{k}^{+}$ averaged over all generated tokens.
@@
  • #toks: This is the number of tokens generated from the expanded vocabulary $V_{k}^{+}\setminus V_{k}$ for a given generation; we report the average over all generations.
@@
As the privacy budget increases, the clipping norm $C$ also increases. We observe that the larger the expansion of the top-$k$ threshold ($2C/B$), the larger the effective-$k$ ($|V_{k}^{+}|$) of Top-$k+$ sampling. Further, as reported in §5, we also observe that the number of tokens sampled from the expanded set is very low, supporting the claim that $V_{k}^{+}$ is a very tight superset of $V_{k}$. The number of sampled tokens also decreases as $k$ increases. Despite a larger increase in effective $k$ due to the expansion of the top-$k$ vocabulary, the probability of selection of tokens decreases sharply with rank since the increased expansion is offset by lower probabilities of selection.
$k=10$ $k=50$ $k=100$ $k=500$ $\bm{\varepsilon}$ $\bm{C}$ Effective ${\bm{k}}$toks Effective ${\bm{k}}$toks Effective ${\bm{k}}$toks Effective ${\bm{k}}$toks $\bm{1.0}$ $0.08$ $10.17$ $6.05$ $50.69$ $1.48$ $101.07$ $0.91$ $506.36$ $0.68$ $\bm{3.0}$ $0.23$ $10.52$ $7.80$ $52.56$ $2.68$ $104.98$ $2.09$ $524.11$ $1.85$ $\bm{5.0}$ $0.36$ $10.52$ $7.77$ $53.09$ $2.87$ $106.82$ $2.62$ $534.65$ $2.50$ $\bm{10.0}$ $0.66$ $11.24$ $10.85$ $56.56$ $4.92$ $113.34$ $4.42$ $568.10$ $4.59$
@@
Table 12: Statistics for Top-$k+$ sampling step of InvisibleInk for $1000$ synthetic text samples generated for the MIMIC dataset using a TinyLLaMA 1B model, batch size of $B=7$ and sampling temperature $\tau=1.2$. We report the size of $V_{k}^{+}$ (effective $k$) and the number of tokens sampled from the expanded vocabulary $V_{k}^{+}\setminus V_{k}$ (#toks) averaged over all generations.
E.3 Ablation Studies: Effect of Temperature and Clip Norm
InvisibleInk has the following key hyperparameters: the batch size $B$ (which is determined by the compute budget), temperature $\tau$, clip norm $C$, and top-$k$ parameter. §5 shows the effect of varying the batch size $B$ (Fig. 4) the top-$k$ parameter (Fig. 6). Here, we study the effect of the other two parameters: temperature $\tau$ and clip norm $C$.
As per the heuristics for hyperparameter selection proposed in §5, we tune across various temperatures and calculate the clip norm $C$ for the fixed privacy budget and sampling temperature $\tau$. The effects of varying the temperature and clip norm are thus coupled. We first report the performance of InvisibleInk for various temperatures; see Fig. 8.
Figure 8: Variation of utility (measured using a variety of metrics) with sampling temperature, for the full-vocabulary variant of InvisibleInk ($k=|V|$), reported for 1000 synthetic generations for the MIMIC dataset using a TinyLLaMA 1B model. Temperature is varied from $0.8-1.1$ for a fixed privacy budget $\varepsilon=10$. We observe that selecting a higher temperature consistently gives better performance.
We now decouple the effect of clipping by removing the constraint on privacy budget and instead varying $C$ freely for a given fixed temperature (see Fig. 9). We also report the variation of calculated clip norm for InvisibleInk and Amin et al. [12]’s method with batch size $B$ for various privacy budgets. As discussed in Fig. 3, the clipping norms required for high-utility generation using InvisibleInk can be achieved with much lower batch sizes as compared to Amin et al. [12]’s method.
Figure 9: Effect of clipping norm. Left two: Variation of calculated $C$ for InvisibleInk and Amin et al. [12]’s method with batch size $B$ for various privacy budgets. The latter needs much larger batch sizes to give comparable clip norms for a given privacy budget. Right two: Variation of utility with $C$ for 1000 synthetic generations for the MIMIC dataset using a TinyLLaMA 1B model using InvisibleInk ($k=|V|$). Temperature ($\tau=1.2$) and batch size ($B=7$) are fixed while we vary the clip norm. Each setting of the clip norm implies a DP parameter as given by Theorem 2. In particular, it varies from $\varepsilon=1.3$ for $C=0.1$ to $\varepsilon=17.6$ for $C=1.0$.
E.4 Comparison of Adjacency Notions for AdaPMixED
As discussed in §B.3, we use an implementation of AdaPMixED that employs the add-or-remove notion of adjacency, as in [18], in the main paper instead of the replace-by-null notion of adjacency used by InvisibleInk. A stricter comparison is offered by using replace-by-null adjacency for AdaPMixED; see Tab. 13. Since the qualitative trends do not change, and InvisibleInk outperforms AdaPMixED irrespective of the notion of adjacency used, the results presented in the main paper use the add-or-remove adjacency for AdaPMixED as proposed in the original paper.
Add-or-remove adjacency Replace-by-null adjacency $\bm{\varepsilon}$ $\bm{B=8}$ $\bm{B=32}$ $\bm{B=128}$ $\bm{B=8}$ $\bm{B=32}$ $\bm{B=128}$ 1.0 $58.68$ $58.71$ $58.51$ $58.14$ $57.64$ $59.51$ 3.0 $58.29$ $58.81$ $61.50$ $58.30$ $57.97$ TLE 5.0 $60.24$ $58.72$ TLE $58.26$ $59.32$ TLE 10.0 $60.23$ $59.23$ TLE $59.18$ $59.61$ TLE
@@ E.4 Comparison of Adjacency Notions for AdaPMixED
Appendix F Example Generations
See Tab. 14 for some example generations from InvisibleInk.
Dataset Examples of Generated Synthetic Text
Yelp "We chose a cozy restaurant that delivered impeccable service, excellent fare, and even better music to our table." (ReviewStars: 4.8 out of 5.0) Naturally, we walked in to find a spacious and serene lobby. Our server was welcoming, and she immediately recommended a seasonal cocktail and the restaurant’s signature dish, Chicken Caesar. The servers were friendly, knowledgeable, and efficient which contributed greatly to our experience. Tucking into each other’s orders, we feasted our eyes on sophisticated appetizers like Caprese, Grilled Salmon, and Grilled Caesar Wings, accompanied by the juiciest and most lip-smacking chicken you’ll ever taste, alongside a delicate glass of Pinot Noir to top it off. The food arrived quickly without forgetting the music
"Although the staff may not be able to befriend every customer, their attentiveness and positive attitude really go a long way. Our recent visit to a new coffee shop led by a chatbot could have gone further in helping patrons’ preferences, with many potential avenues to cater the atmosphere. From a personal observation, however, I believe that adding more personalities to the virtual assistants could cater to a wider-range of needs while engaging customers. Nevertheless, the chatbot is smart and proficient in its tasks, and we enjoyed getting tips about the local coffee culture from their knowledgeable baristas. Although we were a bit inconveniently served at our preferred location for a quick lunchtime fix, we were glad to taste some delicious homemade sushi on the go. Overall, our overall experience was positive with a bit more effort and personalization in terms of addressing specific preferences could lead to repeat vis
TAB PROCEDURE: The case originated in an application to the European Court for Human Rights ("ECHR") filed by the applicant, "John Doe", on this date requesting in particular: - Reasonable time limit for decision on merits of case - Right to information about case proceedings - Right to adequate legal assistance and representation - Right to freedom of assembly and demonstration - Non-compensatory, and not exceptional, measures such as compensation or compensatory or punitive payments - Overseas travel and dissemination of judgment to legal consultants abroad - Enforceability and enforcement of certain judgments and orders - Language of the proceedings, in particular oral and written court proceedings - Procedure for the setting aside of judgments based on Article 6 and Article 13 Annex 1: Applicant’s Criminal Records The applicant had been convicted in absentia by a Russian Court on June 30, 2018, of abuse of power, which, for Russian law purposes, constitutes an example of "organized crime" and is punishable by imprisonment with strict penalties. This conviction constitutes a clear and imminent threat to John Doe’s physical or mental survival. During questioning by the Russian Investigative Committee, John Doe, while in custody, showed "signs of significant deterioration" and "delirium," indicating he was exposed to physical and psychological trauma. Based on these findings, the ECHR held that the applicant was in dire need of life-saving treatment while detained, and that his arrest constitutes a violation of Article 6 of the Convention because he is unable to obtain safe legal representation in his native language, while being held as a material witness in Russian courts. On his current applications, the ECHR: (1) Determined that due to his language barrier and mental instability, it is imperative for him to be allowed to avail himself of fair procedures whereby adequate legal assistance would be provided; (2) Rejected the respondents’ claim based on

### v2

@@ D.6 Confidence Intervals
where $\bar{x}$ is the sample mean, $s$ is the sample standard deviation and $m$ is the number of samples. Here, $Z_{\alpha/2}$ is the $(1-\alpha/2)$ percentile of the standard normal distribution. These are Wald confidence intervals, given under the assumption that error is unbiased and normally distributed. For a $99\%$ confidence interval, we have $Z_{\alpha/2}=2.576$ and for a $95\%$ confidence interval, we have $Z_{\alpha/2}=1.96$.
@@ D.7 Computational Resources
[…]
Evaluation Time. The approximate wall clock runtimes for synthetic text generation using various methods are given below. The runtime varies with the size of the model and batch size used, in particular with the maximum number of parallel LLM inferences that may be run on the GPU. For our experiments using TinyLLaMA, a maximum of $min(B+1,16)$ inferences were run in parallel on 1 GPU. The runtimes also scale (approximately) linearly with the length of individual generated samples and the number of samples generated. All numbers reported below are for $1000$ generations of $500$ token synthetic samples using MIMIC as the reference dataset.
@@
  • AugPE [41]: $\approx 18$ hours for 77 variations per iteration and $10$ iterations
@@ Appendix E Additional Experimental Results
  • Additional results comparing InvisibleInk and AugPE [41] on the MIMIC dataset.
  • Additional experiments on the downstream utility of data generated from the Yelp dataset.
@@
  • Additional results for synthetic generations for the MIMIC dataset using LLaMA3 8B.
  • Additional experiments designed to highlight the effect of sampling temperature $\tau$, clipping norm $C$ and top-$k$ sampling parameter on the quality of generation.
@@ E.1 Comparison with AugPE
Fig.˜8 gives additional experimental results comparing the utility vs computational cost of InvisibleInk and AugPE for various privacy budgets. The trends agree with those in Fig.˜5; see §5.
Figure 8: Utility vs Compute plots (avg. over 3 runs) for InvisibleInk and AugPE for varying privacy budgets $\varepsilon$ for 1000 synthetic texts generated for the MIMIC dataset. We compare utility using a variety of metrics — InvisibleInk outperforms AugPE across all settings and evaluation metrics. Wall-clock run time is used to measure the computational cost. We report results for $B+1=4,8,16,32$ next-token inference calls per generated token and $T_{\mathsf{AugPE}}=1,3,5$ for InvisibleInk and AugPE, respectively.
@@
Table 11: Random sample generation and Variation prompts for MIMIC adapted in the style of AugPE; The candidate synthetic generations at the current iteration are inserted in the space denoted by “<insert candidate synthetic text>” and information about the style in which the LLM should rephrase the text is inserted in the space denoted by “<insert rephrase style>”.
@@
The number of variations of the candidate synthetic texts generated in every iteration of AugPE (the parameter $L$ in [41, Alg 1]) is fixed to be $7$. We also limit the total number of sensitive references the algorithm sees to $7\times n$ (where $n$ is the total number of synthetic texts generated). The total number of full-sequence LLM generations per synthetic text is given by $8\times T_{\mathsf{AugPE}}$, where $T_{\mathsf{AugPE}}$ is the number of iterations, and varies from 8808-80. In contrast, we disadvantage our model by limiting the number of effective full-sequence LLM generations to $B+1$, ie, 4324-32. Further, we use the official implementation of AugPE released by Xie et al. [41] [^13] Available at https://github.com/AI-secure/aug-pe and adapt it for generating synthetic samples from MIMIC with minimal changes. In particular, we preserve the same style of random sample generation and text variation prompts used for other datasets; see Tab.˜11.
E.2 Top-$k$ Selection Statistics
@@
Some additional statistics of selection from the expanded vocabulary described by the Top-$k+$ step of InvisibleInk are presented in Tab.˜12. In particular, we show the following:
  • $k_{\text{eff}}$: Effective-$k$, which is the size of the vocabulary $V_{k}^{+}$
  • $\sigma_{k}$: standard deviation of $k_{\text{eff}}=|V_{k}^{+}|$
@@
  • #toks: number of tokens generated from the expanded vocabulary $V_{k}^{+}\setminus V_{k}$ per generation
@@
For each metric, we report the average over all generations or generated tokens for a setting. As the privacy budget increases, the clipping norm $C$ also increases. We observe that larger expansions of the top-$k$ threshold ($2C/B$), lead to larger effective-$k$ ($k_{\text{eff}}=|V_{k}^{+}|$) of Top-$k+$ sampling. Further, as reported in §5, we also observe that the number of tokens sampled from the expanded set is very low (typically $\leq 10$), supporting the claim that $V_{k}^{+}$ is a very tight superset of $V_{k}$. Further, the standard deviation of $k_{\text{eff}}$ scales roughly with $\sqrt{k_{\text{eff}}-k}$, i.e., the square root of the size of the expansion.
The number of sampled tokens also decreases as $k$ increases. Despite a larger increase in effective $k$ due to the expansion of the top-$k$ vocabulary, the probability of selection of tokens decreases sharply with rank since the increased expansion is offset by lower probabilities of selection.
$k=10$ $k=50$ $k=100$ $k=500$ $\bm{\varepsilon}$ $\bm{C}$ $\bm{k_{\text{eff}}}$ $\bm{\sigma_{k}}$ #toks $\bm{k_{\text{eff}}}$ $\bm{\sigma_{k}}$ #toks $\bm{k_{\text{eff}}}$ $\bm{\sigma_{k}}$ #toks $\bm{k_{\text{eff}}}$ $\bm{\sigma_{k}}$ #toks $\bm{1.0}$ $0.08$ $10.17$ $0.43$ $6.14$ $50.69$ $0.97$ $1.50$ $101.07$ $1.28$ $0.91$ $506.43$ $3.95$ $0.73$ $\bm{3.0}$ $0.23$ $10.52$ $0.76$ $7.71$ $52.57$ $1.77$ $2.62$ $104.98$ $2.54$ $2.04$ $524.04$ $7.11$ $1.81$ $\bm{5.0}$ $0.36$ $10.52$ $0.77$ $7.77$ $53.09$ $2.00$ $2.89$ $106.80$ $3.02$ $2.51$ $534.82$ $9.31$ $2.60$ $\bm{10.0}$ $0.66$ $11.25$ $1.25$ $11.15$ $56.57$ $3.04$ $5.04$ $113.35$ $4.55$ $4.45$ $568.28$ $16.33$ $4.77$
@@
Table 12: Statistics for Top-$k+$ sampling step of InvisibleInk for $1000$ synthetic text samples generated for the MIMIC dataset using a TinyLLaMA 1B model, batch size of $B=7$, and sampling temperature $\tau=1.2$. We report the size of $V_{k}^{+}$ ($k_{\text{eff}}$), the standard deviation in $k_{\text{eff}}$ ($\sigma_{k}$), and the number of tokens sampled from the expanded vocabulary $V_{k}^{+}\setminus V_{k}$ (#toks); all metrics are reported averaged over all generations or generated tokens.
E.3 Scaling InvisibleInk to Larger Models: Results on LLaMA3 8B
The utility scores (MAUVE) for $1000$ synthetic texts generated using a LLaMA3 8B model are presented in Tab.˜13. The results agree qualitatively with trends observed for smaller models: intermediate values of $k\approx 100$ to $1000$ consistently perform better than very small or very large $k$.
$\bm{\varepsilon}$ $\bm{k{=}10}$ $\bm{k{=}50}$ $\bm{k{=}100}$ $\bm{k{=}500}$ $\bm{k{=}1000}$ $\bm{k{=}5000}$
@@ E.3 Scaling InvisibleInk to Larger Models: Results on LLaMA3 8B
[…]
Table 13: MAUVE ($\%$) scores, reported with mean and $95\%$ confidence intervals (see §D.6) over 33 runs, for $1000$ synthetic text generations using LLaMA3 8B for the MIMIC dataset with varying $\varepsilon$ and $k$ for InvisibleInk with $\tau=1.2$ and $B=7$. The best scores for every privacy budget are highlighted.
We note that generation from the LLaMA3 8B at such small batch sizes ($B=7$) is completely infeasible for the other baselines we consider, such as AdaPMixED [19] and Amin et al. [13]’s method. As we use larger (in terms of number of parameters) LLMs, the computational cost in terms of both GPU memory and computational time increases sharply. Using batch sizes >50>50, as proposed by the baselines [19, 13] is unreasonable with large models. This highlights the scalability of InvisibleInk to larger models, and by extension, its suitability for high-quality text generation.
However, we note that the overall MAUVE scores of the generated texts falls compared to generations from smaller models like TinyLLaMA and LLaMA3.2 1B. We attribute this to the strong alignment of the LLaMA3 8B model against generating (fake) medical records or divulging patient data. We further note that despite our detailed prompts (see Tab.˜5), the output private text generated by the model contains text other than the synthetic text to be generated, usually in the form of disclaimers of the form: “Here is a fake patient discharge summary”. Sanitizing the output private text, post-generation, by removing this disclaimer leads to $\approx 1-2\%$ increase in MAUVE scores across all settings.
While our primary experiments (reported in Fig.˜4 and Tab.˜2) show that InvisibleInk performs well, even when generating out-of-domain text, strong alignment of LLMs impacts generation adversely. We leave a full investigation of this phenomenon to future work.
E.4 Ablation Studies: Effect of Temperature, Clip Norm, and Top-$k$ parameter
InvisibleInk has the following key hyperparameters: the batch size $B$ (which is determined by the compute budget), temperature $\tau$, clip norm $C$, and top-$k$ parameter. §5 shows the effect of varying the batch size $B$ (Fig.˜4) the top-$k$ parameter (Fig.˜6).
Figure 9: Variation of utility (measured using a variety of metrics) with sampling temperature, for the full-vocabulary variant of InvisibleInk ($k=|V|$), reported for 1000 synthetic generations, averaged over 33 runs, for the MIMIC dataset using a TinyLLaMA 1B model. Temperature is varied from $0.8$ to $1.1$ for a fixed privacy budget $\varepsilon=5$. We observe that selecting a temperature of $\tau\approx 1.0-1.1$ consistently gives better performance. When coupled with Top-$k+$ sampling for $k\approx 50-100$ (chosen based on empirical observations in Fig.˜2), we observe that the best performance is obtained for $\tau=1.1$. We report these results in the main paper; in Fig.˜4.
As per the heuristics for hyperparameter selection proposed in §5, we tune across various temperatures and calculate the clip norm $C$ for the fixed privacy budget and sampling temperature $\tau$. The effects of varying the temperature and clip norm are thus coupled. We first report the performance of InvisibleInk for various temperatures and top-$k$ parameters. We observe from Fig.˜9 & Fig.˜10 that the optimal choice of sampling temperature and top-$k$ parameter is $\tau=1.1$ and $k=100$.
Figure 10: Variation of utility (measured using MAUVE scores) with sampling temperature and top-$k$ parameter, reported for 1000 synthetic generations, over one run, for the MIMIC dataset using a TinyLLaMA 1B model. Temperature is varied from $1.0$ to $1.2$ for various privacy budgets. We observe that selecting a temperature of $\tau\approx 1.1$ consistently gives the best performance when coupled with Top-$k+$ sampling for $k\approx 50-100$ (chosen based on empirical observations in Fig.˜2). We report these results in the main paper; in Fig.˜4.
We now decouple the effect of clipping by removing the constraint on privacy budget and instead varying $C$ freely for a given fixed temperature (see Fig.˜11). We also report the variation of calculated clip norm for InvisibleInk and Amin et al. [13]’s method with batch size $B$ for various privacy budgets. As discussed in Fig.˜3, the clipping norms required for high-utility generation using InvisibleInk can be achieved with much lower batch sizes as compared to Amin et al. [13].
Figure 11: Effect of clipping norm. Left two: Variation of calculated $C$ for InvisibleInk and Amin et al. [13]’s method with batch size $B$ for various privacy budgets. The latter needs much larger batch sizes to give comparable clip norms for a given privacy budget. Right two: Variation of utility with $C$ for 1000 synthetic generations for the MIMIC dataset using a TinyLLaMA 1B model using InvisibleInk ($k=|V|$). Temperature ($\tau=1.2$) and batch size ($B=7$) are fixed while we vary the clip norm. Each setting of the clip norm implies a DP parameter as given by Theorem˜2. In particular, it varies from $\varepsilon=1.3$ for $C=0.1$ to $\varepsilon=17.6$ for $C=1.0$.
E.5 Comparison of Adjacency Notions for AdaPMixED
As discussed in §B.3, we use an implementation of AdaPMixED that employs the replace-by-null notion of adjacency, as in InvisibleInk, in the main paper instead of the add-or-remove notion of adjacency used by Flemings et al. [19] as it allows for an apples-to-apples comparison.
Add-or-remove adjacency Replace-by-null adjacency $\bm{\varepsilon}$ $\bm{B=8}$ $\bm{B=32}$ $\bm{B=128}$ $\bm{B=8}$ $\bm{B=32}$ $\bm{B=128}$ 1.0 $58.68$ $58.71$ $58.51$ $58.14$ $57.64$ $59.51$ 3.0 $58.29$ $58.81$ $61.50$ $58.30$ $57.97$ $58.80$ 5.0 $60.24$ $58.72$ TLE $58.26$ $59.32$ TLE 10.0 $60.23$ $59.23$ TLE $59.18$ $59.61$ TLE
Table 14: MAUVE scores ($\%$) for $1000$ synthetic generations (over one run only) for the MIMIC dataset using TinyLLaMA for different privacy budgets ($\varepsilon$) and batch sizes ($B$), comparing two adjacency notions: Add-or-remove and Replace-by-null for AdaPMixED. TLE denotes runs where a wall-clock time limit of 24 hrs was exceeded. Top-performing configurations per $\varepsilon$ tend to vary with adjacency notion and batch size; highlighting the sensitivity of MAUVE scores to both.
To do so, we modify the data-dependent privacy accounting step (defined in [19, Eqn. 3]), which accounts for the maximum symmetric $\alpha$-Rényi Divergence between probability distributions corresponding to the reference set and all its possible neighbours under add-or-remove adjacency.
We modify this step to instead account for the maximum $\alpha$-Rényi Divergence between the probability distribution corresponding to the current reference set and all possible neighbours using the replace-by-null adjacency notion. In contrast to the add-or-remove notion, where removing a particular reference zeroed out its contributing probabilities [19, Alg 1, Line 15], in our modified algorithm, replacing a particular reference by “null" (empty string $\emptyset$; see Definition˜5) means that the contributing probabilities default to that of the public distribution. Mathematically, we modify [19, Alg 1, Line 15] to be (borrowing all notation from [19]):
$p_{-i}({\bm{x}})=\frac{1}{N}\left(\sum_{j\neq i}\bar{p}_{j}({\bm{x}})+p_{\mathsf{pub}}({\bm{x}})\right)$
We note that this lends an advantage to the replace-by-null implementation of AdaPMixED, since the per-token privacy accounted for must be slightly lesser. However, the quantitative results, presented in Tab.˜14, fail to reflect this meaningfully, since the generated sequences are very small. Regardless, InvisibleInk outperforms AdaPMixED significantly irrespective of the notion of adjacency used.
Appendix F Software Package
We illustrate the use of the accompanying software package, available on GitHub, [^14] https://github.com/cerai-iitm/invisibleink and installable via pip, [^15] https://pypi.org/project/invink/ as pip install invink.
1import invink
2from datasets import load_dataset
3
4# Load the TAB dataset for this example.
5data = load_dataset("mattmdjaga/text-anonymization-benchmark-train")
6
7# Reference texts given as a list of strings.
8ref_txts = data["train"]["text"]
9
10# Use smaller models for lower memory footprint.
11model = "google/gemma-3-4b-it"
12
13# Description should not contain privacy-sensitive information.
14dataset_desc = "The dataset comprises English-language court cases from the European Court of Human Rights (ECHR)."
15
16# Generate 10 text samples with (10, 1e-5)-DP
17output = invink.generate(ref_txts, model, num=10, epsilon=10,
18 batch_size=8, max_toks=500, dataset_desc=dataset_desc)
19
20# list of private synthetic text strings
21print(output.texts)
Listing 1: Generate private synthetic text using InvisibleInk
[…]
Appendix G Example Generations
See Tab.˜15 for some example generations from InvisibleInk.
Dataset Examples of Generated Synthetic Text Yelp "We chose a cozy restaurant that delivered impeccable service, excellent fare, and even better music to our table." (ReviewStars: 4.8 out of 5.0) Naturally, we walked in to find a spacious and serene lobby. Our server was welcoming, and she immediately recommended a seasonal cocktail and the restaurant’s signature dish, Chicken Caesar. The servers were friendly, knowledgeable, and efficient which contributed greatly to our experience. Tucking into each other’s orders, we feasted our eyes on sophisticated appetizers like Caprese, Grilled Salmon, and Grilled Caesar Wings, accompanied by the juiciest and most lip-smacking chicken you’ll ever taste, alongside a delicate glass of Pinot Noir to top it off. The food arrived quickly without forgetting the music "Although the staff may not be able to befriend every customer, their attentiveness and positive attitude really go a long way. Our recent visit to a new coffee shop led by a chatbot could have gone further in helping patrons’ preferences, with many potential avenues to cater the atmosphere. From a personal observation, however, I believe that adding more personalities to the virtual assistants could cater to a wider-range of needs while engaging customers. Nevertheless, the chatbot is smart and proficient in its tasks, and we enjoyed getting tips about the local coffee culture from their knowledgeable baristas. Although we were a bit inconveniently served at our preferred location for a quick lunchtime fix, we were glad to taste some delicious homemade sushi on the go. Overall, our overall experience was positive with a bit more effort and personalization in terms of addressing specific preferences could lead to repeat vis TAB PROCEDŽURE: The case originated in an Application for Judicial Review filed by Ms. Smith, petitioner(s), against the Bulgarian State from January 8th of the year 2019. The Respondent for this jurisdictional issue is the Ministry of the Interior of the Bulgarian Republic and the Bulgarian National Antiterrorism Bureau. The alleged event occurred at the intersection of a public place known as Sidi Boutha Crescent or Vezhub Khayri Crescent in the city of Sofia, Bulgaria that caused or contributed significantly and directly to plaintiff Smith’s physical injury causing death pursuant to Section 4, Article 3 of the Bulgarian Medical Registration Act. - The procedural steps involved in the application included the preparing of the written petition as well as its formality submitted to the Court for review. - Upon review of the documents, the Registrar of the Court sent the petition to the Bulgarian Ministry of the Interior as a legal source and required from the said organization to verify and respond in court about the availability of sufficient grounds for such judicial review against the Bulgarian state. - Following the review process, the request for verification made by plaintiff Smith’s attorney was submitted before the Ministry of the Interior of the Bulgarian Republic. - The Bulgarian Minister of the Interior received the request as per the requirement and made a decision within one month from the submission date based on the relevant document provisions as provided by law. - After this, the Ministry communicated to the plaintiff regarding the reply received from the National Antiterrorism Bureau, and the said communication stated that the respondent is satisfied with having received the reply and that the Bulgarian state is willing to conduct a judicial investigation of the event from the incident that caused or contributed significantly to the death of the petitioning defendant Smith. Respectfully, as per the international human rights instrument cited in the previous reference to the jurisdiction as mentioned in the jurisdictional challenge brought in favor of Ms. Smith, all facts and circumstances of the criminal matter that led to the loss of a life of a Bulgarian citizen constitutes a violation of her right to physical integrity protected by Article 2 (Prohibition of T